A 5.00 mL saturated sample of borax solution was collected at 43.8 degrees Celsius. It required 16.39 mL of 0.2503 M HCl to titrate. Calculate the Ksp for this sample.
1
Expert's answer
2019-03-27T05:34:05-0400

n(HCl) = C(HCl) × V(HCl) = 0.2503M × 0.01639L = 0.0041mol.

  Xmol                0.0041mol

Na2B4O7×10H2O + 2HCl = 2NaCl + 4H3BO3 + 5H2O

 1 mol                    2mol


nborax = Xmol = 0.0041/2 mol = 0.00205mol.

Cborax = nborax/Vborax = 0.00205mol/0.005L = 0.41M.

 0.41M                   0.82M 0.41M

Na2B4O7×10H2O  ⇄ 2Na+ + B4O72- +10H2O

    1M                        2M       1M

Ksp borax = [Na+]2 × [B4O72-] = 0.822 × 0.41 = 0.2757 ≈ 0.28.